# Spatial structure facilitates evolutionary rescue by cost-free drug resistance

### v1

@@ Results
To understand in more detail the role of spatial structure on resistance spread, let us study the system composition before drug is added.
Figure 3A shows the time evolution of the average number of R mutants present in the whole structured population. We observe that it is not impacted by population structure. Whatever the migration rate, the number of R individuals grows as mutants fix in some replicates of our stochastic simulations. This occurs either first locally for structured populations, or directly in the whole population for the well-mixed population (see SI Section 3.3), but the average over replicates remains the same. The long-time limit corresponds to the total steady-state population size, $N=K\left(1-g/f_{R}\right)$, as R mutants take over the whole population. This occurs with a timescale of order $\langle t_{af\,W}\rangle$. Figure 3B shows that the variance across replicates of the number of R mutants is strongly impacted by population structure. This is particularly true in the time interval between $\langle t_{af\,F}\rangle$ and $\langle t_{af\,W}\rangle$. This suggests that the impact of population structure on the variance of mutant number is driven by R mutant fixation. Figure 3B further shows that the variance of the number of mutants across replicates is larger when migrations are more frequent. This can be understood qualitatively by comparing a well-mixed population and a structured population with small migration rate. In the first case, mutants fix in one step, while in the second one they fix separately in each deme (see Figure S2). The large variance observed for the well-mixed population comes from the variability across replicates of the appearance time of a successful mutant. For the structured population with small migration rate, this is partly smoothed by lumping together $D$ separate demes with different fixation times. Moreover, the appearance time of a successful mutant in the fastest deme has a smaller variance than that in a well-mixed population (see SI Section 3.3). Here, we focused on mutant fraction in the whole population because we are interested in the overall behavior of the population, but the dynamics at the deme level is described in SI Section 3 and below in Figure 4.
@@
Figure 3: Impact of spatial structure on population composition in the absence of drug. Panel A: The mean across replicates of the total number of R mutants in the population is shown versus time for population structures differing by the migration rate $\gamma$, and for the well-mixed population with same total size (“WMP”). Horizontal dashed line: steady-state total population size. Panel B: Variance across replicates in the total number of R mutants in the population, in the same conditions as in panel A. Vertical dotted line in panels A and B: time when largest variance was obtained in panel B (t$=5\times 10^{5}$). Panel C: Population composition at t$=5\times 10^{5}$, for different migration rates. Results are obtained from $10^{4}$ replicate simulations. Parameter values (in all panels): $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
@@
To go beyond the mean and variance in overall mutant number, in Figure 3C we examine the population composition in the absence of drug, at the time when we obtained the largest variance in Figure 3B (vertical dotted line). We describe the population composition by defining four categories: either there are no R bacteria at all (“No R” - red), or at least one deme comprises between 1 and 9 R bacteria but no deme has more than 9 R (“Small R population” - yellow), or at least one deme comprises more than 9 R individuals, but the R type has not fixed in the whole population (“Big R population” - cyan), or R bacteria have fixed in the whole population (“Only R” - blue). Note that we distinguish “Small R population” from “Big R population” to account for the possibility of stochastic extinction of R lineages, which becomes negligible if at least 10 R bacteria are present in a deme (see SI Section 3). In Figure 3C, we observe that in most replicates, at the time considered, the well-mixed population has either fixed resistance or does not have any mutants. Thus, the number of mutants is either 0 or $(1-g/f_{R})KD$, yielding the large variance mentioned above. As we partition the system spatially and lower the migration rate, the composition of structured populations becomes more mixed, and the fraction of replicates with overall fixation or overall extinction decreases. This results in a smaller variance across replicates, as the number of mutants becomes more homogeneous across them.
@@
An important cause of the composition difference between structured and well-mixed populations is the possibility of local fixation of R mutants. What is the fixation dynamics at the deme level before we add the drug? In Figure 4, we report how many demes have fixed resistance for different migration rates and at different times, in the form of histograms computed over simulation replicates. Note that for the well-mixed population we report overall fixation. As expected, the number of demes where R mutants have fixed increases over time. Furthermore, we observe that the distribution of demes that have fixed resistance and its dynamics are strongly impacted by spatial structure. For large migration rates (panel A), at all times, most replicates have fixed resistance in either no deme or all demes. This is close to the large-$\gamma$ limit (i.e. the well-mixed population), shown as horizontal lines in panel A. When the migration rate $\gamma$ is decreased (panels B and C), the fraction of realizations featuring an intermediate number of demes that have fixed resistance increases at intermediate times. For the small migration rate $\gamma=10^{-7}$ (panel C), the distributions become close to those observed for $\gamma=0$ (panel D). Thus, small migration rates result in a population with a transient strong heterogeneity across demes. This has a crucial impact on the outcome of drug treatment, and on the survival probability curves in Figure 2. Indeed, the biostatic drug does not affect R mutants, and the presence of at least one deme where R has fixed rescues the population when the drug is applied.
@@
Figure 4: Number of demes that have fixed resistance before drug application. Histograms of the number of demes that have fixed resistance, computed across $10^{4}$ simulation replicates, are shown for different values of the elapsed time t (on different curves with different colors), and of the migration rate $\gamma$ (in different panels). Specifically, for each $i\in[0,10]$, we show the fraction of realizations in which $i$ demes have fixed resistance. The horizontal line markers in panel A show the results for a well-mixed population with same total size (“WMP”). Parameter values: $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
@@
So far, we focused on the impact of spatial structure on the survival probability of the population, and on its composition before the addition of drug. A key conclusion is that local fixation of R mutants allows earlier population survival. Let us now ask what happens after drug addition. Since our biostatic drug prevents S bacteria from dividing while mutations occur upon division, there are no new mutations after drug addition. If R mutants fixed in at least one deme before drug addition, how fast do they spread in all demes of a spatially structured population?
@@
To address this question, let us first calculate the time $\tau_{c\,mig}$ it takes for an R mutant to migrate to a deme without R mutants, starting from a population where resistance has fixed in $k\geq 1$ demes out of $D$. Then, there are $Nk$ mutant individuals, where $N$ is the steady-state deme size, which can migrate to each of the $D-k$ uncolonized demes at a per capita rate $\gamma$. However, once an R mutant arrives in a new deme, its lineage may go stochastically extinct. This happens with probability $1-g/f_{R}$, assuming that the deme does not contain other individuals (see derivation in SI Section 1.5). Thus, we have:
$\tau_{c\,mig}(k)=\frac{1}{\gamma Nk(D-k)}\frac{1}{1-g/f_{R}}=\frac{1}{\gamma Nk(D-k)}\,C\,,$ (3)
@@
where $C=1/(1-g/f_{R})$ captures the effect of stochastic extinctions. In Figure 5A, we compare the analytical prediction in Equation (3) to our simulation results, obtaining a good agreement. We also show the impact of stochastic extinctions by setting $C=1$ in the analytical predictions: their effect is not negligible here (of order 10%, and they would be larger if $g$ was closer to $f_{R}$). The time $\tau_{c\,mig}$ for R mutants to colonize the next deme features a minimum when half of the demes are mutant. A larger proportion of migration events then leads to colonization. Indeed, more mutant demes yield more mutants that may migrate, but fewer wild-type demes where they can fix, leading to a trade-off.
@@
Equation (3) neglects the presence of S bacteria in the population, which would affect the stochastic extinction of R lineages. This is acceptable provided that the extinction of S bacteria upon drug addition is fast enough. The decay time $\tau_{S}$ of a well-mixed population comprising $N$ S bacteria upon drug addition reads [marrec_resist_2020]:
$\tau_{S}=\frac{1}{g}\sum_{i=1}^{N}\frac{1}{i}\,.$ (4)
For the migration rate considered in Figure 5A, we checked that $\tau_{S}$ is indeed negligible with respect to $\tau_{c\,mig}$.
@@
Figure 5: Colonization of the whole population by R mutants. Panel A: The time $\tau_{c\,mig}$ to colonization of the next deme by R mutants through migrations is shown versus the number $k$ of demes where R mutants have already fixed. Migration rate: $\gamma=10^{-6}$. We compare simulation data from $5\times 10^{3}$ realizations with the theoretical prediction from Equation (3), accounting for stochastic extinctions (“Theory”) or not (“Theory, $C=1$”). Error bars represent 95% CI. Panel B: The average time to complete colonization (i.e., when population size first becomes at least equal to $0.9K$ in each separate deme) is shown versus the drug addition time $t_{add}$ for structured populations with different migration rates, and for the well-mixed population with same total size (“WMP”). Out of $10^{4}$ realizations, we restrict to simulations where the population survived treatment. Solid black line: $y=x$ diagonal. Horizontal lines: $t_{c\,tot}(\gamma)$ for the two lowest migration rates $\gamma$. Vertical lines: average local fixation time of R mutants in the fastest deme ($\langle t_{af\,F}\rangle$), in a well-mixed population ($\langle t_{af\,W}\rangle$), and in the slowest deme ($\langle t_{af\,S}\rangle$). Confidence intervals are too small to be visible in this logarithmic scale. Parameter values for both panels: $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
So far, we focused on one step of the spread of R mutants. How long does it take in practice for R mutants to colonize the whole structured population? In Figure 5B, we show the total colonization time versus drug addition time for different migration rates. We observe different regimes in these curves depending on the drug addition time $t_{add}$.
First, if the drug is added after a short time, at most one deme in the population will comprise R mutants or will have fixed them. This should happen if the drug addition time $t_{add}$ is shorter than the average appearance time $\langle t_{af\,F}\rangle$ of a locally successful mutant in the fastest deme. How long does it then take for R mutants to colonize the whole population? To obtain this total colonization time $t_{c\,tot}$, we can sum $\tau_{c\,mig}$ over the $D-1$ steps needed to sequentially colonize the structured system. It is thus given by:
$t_{c\,tot}=\sum_{k=1}^{D-1}\tau_{c\,mig}(k)=\frac{2(\Gamma+\psi(D))}{D\gamma N}\frac{1}{1-g/f_{R}}\,,$ (5)
@@
where $\Gamma$ is the Euler gamma constant, while $\psi$ is the digamma function. In Figure 5B, we report the overall colonization time of the population by R mutants, defined as the time $t_{\textrm{add}}$ when drug is added plus the time needed for colonization after drug addition. We observe that for small addition times satisfying both $t_{add}\ll\langle t_{af\,F}\rangle$ and $t_{add}\ll t_{c\,tot}$, $t_{c\,tot}$ indeed describes well the colonization time. Indeed, it is then unlikely that mutants are present in more than one deme when drug is added, and the addition time has a negligible contribution. It leads to a plateau whose value is governed by the migration rate. Note that in Figure 5B, this plateau is only observed for small migration rates ($\gamma\lesssim 10^{-5}$) such that the condition $t_{add}\ll t_{c\,tot}$ is satisfied for the smallest $t_{add}$ values considered. For larger values of $t_{add}$ and larger migration rates, the timescale $t_{add}$ dominates the overall colonization time of the population by R mutants, as we have $t_{add}>t_{c\,tot}$. In these cases, mutant spread can be considered fast when drug is added (see Figure 5B). Finally, if $t_{add}$ is further increased, overall colonization of the population by R mutants may happen before drug addition. Thus, the overall colonization time becomes smaller than $t_{add}$ and tends toward a new plateau. Note that in this regime, colonization can occur before drug addition through migrations or new mutations, followed by fixations. The time it takes for a locally successful mutant to appear in the slowest deme can be expressed as $\langle t_{af\,S}\rangle=\langle t_{af\,W}\rangle\sum_{i=1}^{D}1/i$ (see SI Section 1.2). It is beyond this time that we expect a plateau, even for small migration rates. This is indeed what is observed in Figure 5B.
Extension to cases where resistance carries a cost or a benefit
So far, we considered that the resistant mutant is neutral. However, antibiotic resistance often carries a cost in the absence of drug [Andersson99, Andersson10]. Besides, if antibiotic is already present at relatively low doses in the environment before the addition of drug, it can induce a benefit of resistance before drug is added. Let us briefly consider these cases.
First, our results extend to effectively neutral mutants having a fitness cost $\delta$ such that $N\delta\ll 1$. The impact of structure should in fact be even stronger if $N\delta\ll 1$ but $ND\delta\gg 1$ because fixation is then strongly suppressed in a well-mixed population of size $ND$. Second, the case where the cost of resistance $\delta$ is stronger ($N\delta\gg 1$) is discussed in SI Section 3.1. While spatial structure should have a strong effect here, this will only matter for the very long times associated to the fixation of deleterious mutants. Third, the effect of spatial structure we evidenced here for neutral mutants does not extend to the case where resistant mutants are beneficial, see SI Section 3.1.
Extension to an inoculum that contains R mutants
So far, we focused on a sensitive inoculum and on resistant individuals appearing through mutations. The case where mutants are already present in the inoculum is also interesting. We consider it in SI Section 5. Our main results about the increase in probability of treatment survival granted by a spatial population structure extends to the case where R mutants are present in the inoculum (see also Ref. [kreger_role_2023]). Specifically, we observe that in this case too, the survival probability of the population increases when the mutation rate decreases, see Figure S6.
More complex spatial structures behave similarly to the clique
So far, we considered a minimal model of spatially structured population, known as the clique, where all demes are equivalent and connected to one another by identical migration rates. Do our findings extend to more complex spatial structures? To address this question, we explored two structures with more reduced symmetry.
First, we consider a two-dimensional square lattice with periodic boundary conditions. Here, each deme is directly connected to its four nearest neighbors, thus only allowing for local migrations. We compare the lattice to the clique with the same total outgoing migration rate from a deme. Our simulations yield no statistically significant difference for the survival probability in the square lattice and in the clique (see SI Figure S5). This suggests that making migrations more local does not affect treatment survival.

### v2

@@ Results
To understand in more detail the role of spatial structure on resistance spread, let us study the system composition before drug is added. For simplicity, let us perform this analysis in the case of neutral R mutants, with the same parameter values as in Fig 2A.
Fig 3A shows the time evolution of the total number of R mutants present in a structured population, averaged over many stochastic simulation replicates. Whatever the migration rate, the average number of R individuals grows as mutants fix in some simulation replicates. This occurs either first locally for structured populations, or directly in the whole population for the well-mixed population (see Supplementary Appendix Section 3.3). Despite this difference, the average over replicates is not impacted by population structure. Indeed, for neutral mutants, using Eq 1 and $p_{\textrm{fix}}(N^{*})=1/N^{*}$ gives an average time $\langle t_{af\,W}(N^{*})\rangle=1/(\mu g)$ for a successful mutant to appear in a well-mixed population. As this timescale is independent of population size, it governs the process of R mutants taking over each single deme but also a larger well-mixed population.
@@
Figure 3: Impact of spatial structure on population composition in the absence of drug. Panel A: The mean across replicates of the total number of R mutants in the population is shown versus time for population structures differing by the migration rate $\gamma$, and for the well-mixed population with same total size (“WMP”). Horizontal dashed line: steady-state total population size $K\left(1-g/f_{R}\right)$. Panel B: Variance across replicates in the total number of R mutants in the population, in the same conditions as in panel A. Vertical dotted line in panels A and B: time when largest variance was obtained in panel B ($t=5\times 10^{5}$). Panel C: Population composition at $t=5\times 10^{5}$, for different migration rates. We distinguish four cases. “Only R”: R bacteria have fixed in the whole population; “Large R population”: at least one deme comprises more than 9 R individuals, but the R type has not fixed in the whole population; “Small R population”: at least one deme comprises between 1 and 9 R bacteria, but no deme has more than 9 R; “No R”: there are no R bacteria. We distinguish “Large” and “Small R population” to account for the possibility of stochastic extinction of R lineages, which becomes negligible if $\gtrsim 10$ R bacteria are present in a deme (see Supplementary Appendix Section 3.4). Results are obtained from $10^{4}$ replicate simulations. Parameter values (in all simulations): $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$. Left and right of panel C: Schematics showing typical population states in four replicates (square boxes) for a well-mixed (left) and a fully subdivided population composed of $D=4$ demes (right). Blue: S bacteria; orange: R bacteria.
Fig 3B shows that the variance across replicates of the total number of R mutants in a population is strongly impacted by its structure. This is particularly true in the time interval between $\langle t_{af\,F}(N^{*},D)\rangle$ and $\langle t_{af\,W}(DN^{*})\rangle$ (see Fig 2A). This suggests that the impact of population structure on the variance of mutant number is driven by R mutant fixation. Fig 3B further shows that the variance of the number of mutants across replicates is larger when migrations are more frequent. In a well-mixed population, mutants fix in one step, while in a fully subdivided population, they fix separately in each deme (see Supplementary Appendix Section 3.3). The large variance observed for the well-mixed population comes from the variability across replicates of the appearance time of a successful mutant. For the fully subdivided population, this is smoothed by lumping together $D$ separate demes. Within each deme, the variance across replicates of the number of R mutants is not impacted by spatial structure, if R mutants are neutral, see Supplementary Appendix Section 3.5.
@@
In Fig 3C, we examine in more detail the population composition in the absence of drug, at the time when we obtained the largest variance in Fig 3B (vertical dotted line). We observe that in most replicates, the well-mixed population has either fixed resistance or does not have any R mutants. Thus, the number of mutants is either 0 or $(1-g/f_{R})DK$, yielding the large variance seen in Fig 3B. Typical compositions of the well-mixed population are illustrated schematically on the left side of Fig 3C. As we partition the system spatially and reduce the migration rate, structured populations more frequently include some fully R demes and some fully S ones. Consequently, the fraction of replicates showing overall fixation or overall extinction decreases. This results in a smaller variance across replicates, as the number of mutants becomes more homogeneous across them. Typical compositions of the fully subdivided population are illustrated schematically on the right side of Fig 3C.
@@
An important cause of the composition difference between structured and well-mixed populations is the possibility of local fixation of R mutants. What is the fixation dynamics at the deme level before we add the drug? In Fig 4, we report how many demes have fixed resistance for different migration rates and at different times, in the form of histograms computed over simulation replicates. Note that for the well-mixed population we report overall fixation. As expected, the number of demes where R mutants have fixed increases over time. Furthermore, we observe that the distribution of demes that have fixed resistance and its dynamics are strongly impacted by spatial structure. For large migration rates (panel A), at all times, most replicates have fixed resistance in either no deme or all demes. This is close to the large-$\gamma$ limit (i.e. the well-mixed population), shown as horizontal lines in panel A. When the migration rate $\gamma$ is decreased (panels B and C), the fraction of realizations featuring an intermediate number of demes that have fixed resistance increases at intermediate times. For the small migration rate $\gamma=10^{-7}$ (panel C), the distributions become close to those observed for $\gamma=0$ (panel D). Thus, small migration rates result in a population with a transient strong heterogeneity across demes. This has a crucial impact on the outcome of drug treatment, and on the survival probability curves in Fig 2. Indeed, the biostatic drug does not affect R mutants, and the presence of at least one deme where R has fixed rescues the population when the drug is applied. Population composition is further shown versus time for different migration rates in the Supplementary Appendix Section 3.6.
@@
Figure 4: Number of demes that have fixed resistance before drug application. Histograms of the number of demes that have fixed resistance, computed across $10^{4}$ simulation replicates, are shown for different values of the elapsed time $t$ (different colors), and of the migration rate $\gamma$ (different panels). Specifically, for each $i\in[0,10]$, we show the fraction of realizations in which $i$ demes have fixed resistance. The horizontal line markers in panel A show the results for a well-mixed population with same total size (“WMP”). Parameter values: $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
Spatial structure facilitates rescue by resistance in broad parameter regimes
Mutation rate and population size.
As mentioned in “Models and methods”, we focus on the rare mutation regime $DK\mu\ll 1$. Outside of this regime, R mutants will usually be present in the population, and allow its rescue, independently of spatial structure. Another point where population size could have a role is that it takes longer for a lineage of R mutants destined to fix in a deme to actually fix there if the deme is larger. However, this does not affect our rescue scenario, because as long as the number of R mutants that are present when drug is added is sufficient for stochastic extinction not to occur after the addition of drug, the population will be rescued.
Cost of resistance.
For neutral resistant mutants ($\delta=0$), we saw that a locally successful mutant appears in a structured population $D$ times faster than a successful mutant in a well-mixed population with same total steady-state size $DN^{*}$, where $N^{*}$ is the steady-state size of a deme. This leads to a substantial effect of structure on evolutionary rescue by drug resistance, see Fig 2A. For mutants with a substantial cost of resistance $\delta\gg 1/K$, the mutant fixation probability in a deme is $p_{\textrm{fix}}(N^{*})\approx\delta e^{-N^{*}\delta}$ (see Supplementary Appendix Section 1.2), and thus the time until a successful mutant appears in a well-mixed population grows exponentially with $N^{*}$ (see Eq 1). The effect of spatial structure on rescue by resistance is then even stronger than for neutral mutants, but it exists at exponentially longer timescales (see Supplementary Appendix Section 3.2). Finally, mutants with intermediate cost satisfying $1/(DK)\ll\delta\ll 1/K$ are effectively neutral in demes of size $N^{*}$, but substantially deleterious in a well-mixed population of size $DN^{*}$. This regime yields a strong effect of spatial structure, which extends from the same timescales as in the neutral case to longer ones, as in Fig 2B (note that mutants are slightly more costly there: $\delta=1/K$). Thus, spatial structure facilitates evolutionary rescue by resistance for all costs $\delta\geq 0$, but the effect exists for extremely long timescales if $\delta\gg 1/K$. Because of this, we focus on small costs.
Migration rate.
So far, we mainly considered the extreme cases of the fully subdivided population and of the well-mixed population for our analytical reasoning. With the parameter values chosen in Fig 2, we observe that in practice, the transition between these two extreme cases occurs when the per capita migration rate is increased from $\gamma=10^{-7}$ to $\gamma=10^{-4}$. More generally, what is the relevant range of migration rates where spatial structure facilitates evolutionary rescue by drug resistance? Given the discussion above on the cost of resistance, let us focus on R mutants with cost satisfying $\delta\ll 1/K$. These mutants are neutral or effectively neutral within demes, and the fixation probability of an R (resp. S) individual in a deme of S (resp. R) bacteria is $p_{\textrm{fix}}(N^{*})=1/N^{*}$.
We determined that spatial structure can facilitate evolutionary rescue by drug resistance because it fosters local fixation of R mutants. The demes that fix R mutants act as refugia for R mutants, and allow population survival when drug is added. For this to happen, local fixation of R mutants must be possible, without being perturbed by S bacteria migrating from other demes during the fixation process. The average fixation time of a neutral mutant in a well-mixed population of size $N^{*}$ is $\langle t_{\textrm{fix}}\rangle=(N^{*}-1)/g$, where we used the Moran process result [ewens_mathematical_2004]. Meanwhile, the expected time until an S individual with a lineage destined for local fixation migrates into the deme where R is in the process of taking over is $\langle t_{\textrm{mig}}\rangle=1/[N^{*}\gamma(D-1)p_{\textrm{fix}}(N^{*})]=1/[\gamma(D-1)]$. Thus, for incoming migrations not to strongly perturb local fixations, we need $\langle t_{\textrm{fix}}\rangle\ll\langle t_{\textrm{mig}}\rangle$, i.e.
$\frac{\gamma}{g}\ll\frac{1}{(D-1)(N^{*}-1)}\,.$ (5)
Accordingly, in Fig 2A, spatial structure significantly impacts population survival when Eq 5 holds, i.e. when $\gamma\ll 10^{-4}$, and results become similar to those of a well-mixed population when $\gamma=10^{-4}$. Note that Eq 5 is less stringent than the rare migration regime [slatkin_fixation_1981, hauert_fixation_2014, marrec_toward_2021], which requests a separation of timescales between migration and fixation processes. Here, we took into account the fact that only some migration events may strongly impact deme composition.
Once the fastest deme has fixed resistance, S bacteria may still migrate to that deme and fix there, leading to the extinction of R mutants. The expected time until this happens is again $\langle t_{\textrm{mig}}\rangle=1/[\gamma(D-1)]$. Let us compare this to the average appearance time of a mutant that fixes in the fastest deme, given in Eq 2, i.e. $\langle t_{af\,F}(N^{*},D)\rangle=1/(D\mu g)$ for effectively neutral mutants. If $\langle t_{\textrm{mig}}\rangle\gg\langle t_{af\,F}(N^{*},D)\rangle$, R mutants that locally fixed are expected to survive until other locally successful mutants appear. Then, provided that $t_{\textrm{add}}\gg\langle t_{af\,F}(N^{*},D)\rangle$, R mutants are present when drug is added, leading to rescue by resistance. Thus, spatial structure is expected to facilitate evolutionary rescue by resistance about as much as in a fully subdivided population if $\langle t_{\textrm{mig}}\rangle\gg\langle t_{af\,F}(N^{*},D)\rangle$, i.e.
$\frac{\gamma}{\mu g}\ll\frac{D}{D-1}\approx 1\,.$ (6)
Indeed, in Fig 2A, structured populations with migrations and fully subdivided ones have very similar survival probabilities when $\gamma\ll 10^{-6}$, i.e. when Eq 6 is satisfied. Survival probabilities intermediate between those of well-mixed populations and of fully subdivided populations are observed when Eq 5 is satisfied but not Eq 6. Note that while $\langle t_{\textrm{mig}}\rangle$ is the expected time until the next S individual destined for fixation migrates into the single R deme, R individuals from that deme may migrate to other demes. Taking this into account gives an expected extinction time of R mutants which is $\langle t_{\textrm{mig}}\rangle$ times a prefactor of order $\log(D)$, assuming rare migrations [bitbol_quantifying_2014].
@@
So far, we focused on the impact of spatial structure on the survival probability of the population upon drug addition. Let us now ask what happens after drug addition. With biostatic drug, S bacteria cannot divide or mutate, and their decay leaves some demes empty. If R mutants rescue a spatially structured population because they were present in at least one deme before drug addition, how fast do they then spread in all demes? This question has implications for disease recurrence, where after an initial transient alleviation of symptoms due to the drug-induced decay of S bacteria, the spread of R mutants could lead to a regrowth of the pathogenic bacteria population.
@@
To address this question, let us first calculate the time $\langle t_{\textrm{c mig}}\rangle$ it takes for an R mutant to colonize an empty deme, starting from a population where resistance has fixed in $k\geq 1$ demes out of $D$, while other demes are empty. Then, there are $N^{*}k$ mutant individuals, where $N^{*}$ is the steady-state deme size, which can migrate to each of the $D-k$ empty demes at a per capita rate $\gamma$. However, once an R mutant arrives in an empty deme, its lineage may go stochastically extinct. This happens with probability $1-g/f_{R}$ (see Supplementary Appendix Section 1.1). Thus, we have:
$\langle t_{\textrm{c mig}}(k)\rangle=\frac{1}{\gamma N^{*}k(D-k)(1-g/f_{R})}\,.$ (7)
@@
In Fig 5A, we compare the analytical prediction in Eq 7 to our simulation results, obtaining a good agreement. The time $\langle t_{\textrm{c mig}}\rangle$ for R mutants to colonize the next deme features a minimum when half of the demes are mutant. Indeed, more mutant demes yield more mutants that may migrate, but fewer wild-type demes where they can fix, leading to a trade-off.
@@
Eq 7 neglects the presence of S bacteria in the population, which would affect the stochastic extinction of R lineages. This is acceptable provided that the extinction of S bacteria upon drug addition is fast enough. The decay time upon drug addition of a well-mixed population comprising $N^{*}$ S bacteria is $\tau_{S}=\left(1/g\right)\sum_{i=1}^{N^{*}}1/i$ [marrec_resist_2020]. For the migration rate considered in Fig 5A, $\tau_{S}$ is indeed negligible with respect to $\langle t_{\textrm{c mig}}\rangle$ (see Supplementary Appendix Section 3.7). More generally, the condition $\tau_{S}\ll\langle t_{\textrm{c mig}}(1)\rangle$ yields $\gamma/g\ll 1/[N^{*}(D-1)\sum_{i=1}^{N^{*}}1/i]$, which is holds in most cases where spatial structure favors rescue by resistance (see Eq 5). The condition $\tau_{S}\ll\langle t_{\textrm{c mig}}(N^{*}/2)\rangle$ is however slightly more demanding, as it yields $\gamma/g\ll 4/[N^{*}D^{2}\sum_{i=1}^{N^{*}}1/i]$.
@@
Figure 5: Colonization of the whole population by R mutants. Panel A: The time $\langle t_{\textrm{c mig}}\rangle$ to colonization of the next deme by R mutants through migrations after drug addition is shown versus the number $k$ of demes where R mutants have already fixed. Migration rate: $\gamma=10^{-6}$. We compare simulation results from $5\times 10^{3}$ realizations with the theoretical prediction from Eq 7 (“Theory”). Error bars represent 95% CI. Panel B: The average time to complete colonization by R mutants (i.e., when the number of R mutants first becomes at least equal to $0.9K$ in each separate deme) is shown versus the drug addition time $t_{\textrm{add}}$ for structured populations with different migration rates $\gamma$, and for the well-mixed population with same total size (“WMP”). Out of $10^{4}$ replicate simulations, we restrict to those where the population survived treatment. Horizontal lines: analytical prediction for $\langle t_{\textrm{c tot}}\rangle$ from Eq S36 for the two lowest $\gamma$ values. Vertical lines: average local fixation time of R mutants in the fastest deme ($\langle t_{af\,F}(N^{*},D)\rangle$), in a well-mixed population ($\langle t_{af\,W}(DN^{*})\rangle$), and in the slowest deme ($\langle t_{af\,S}(N^{*},D)\rangle$). The black line $x=y$ separates colonization occurring before drug is added (below the line) and after (above it). Confidence intervals are too small to be visible in this logarithmic scale. Parameter values for both panels: $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
@@
So far, we focused on one step of the spread of R mutants. How long does it take in practice for R mutants to colonize the whole structured population after drug addition? First, if drug is added after a short time $t_{\textrm{add}}<\langle t_{af\,F}(N^{*},D)\rangle$, at most one deme in the population should comprise R mutants or have fixed them. Then, the average time $\langle t_{\textrm{c tot}}\rangle$ it takes for R mutants to colonize the whole population can be obtained by summing the step-by-step colonization times in Eq 7, see Supplementary Appendix Section 3.7. In Fig 5B, we report the overall colonization time of the population by R mutants, defined as the time $t_{\textrm{add}}$ when drug is added plus the time needed for colonization after drug addition. We observe that for small addition times satisfying both $t_{\textrm{add}}\ll\langle t_{af\,F}(N^{*},D)\rangle$ and $t_{\textrm{add}}\ll\langle t_{\textrm{c tot}}\rangle$, the colonization time is indeed well-described by $\langle t_{\textrm{c tot}}\rangle$. It leads to a plateau whose value is governed by the migration rate. Note that in Fig 5B, this plateau is only observed for small migration rates ($\gamma\lesssim 10^{-5}$) such that the condition $t_{\textrm{add}}\ll\langle t_{\textrm{c tot}}\rangle$ is satisfied for the smallest $t_{\textrm{add}}$ values considered. For larger values of $t_{\textrm{add}}$ and larger migration rates, the timescale $t_{\textrm{add}}$ dominates the overall colonization time of the population by R mutants, as $t_{\textrm{add}}>\langle t_{\textrm{c tot}}\rangle$, see Fig 5B. In these cases, mutant spread can be considered fast when drug is added. Finally, if $t_{\textrm{add}}$ is further increased, R mutants may fix in the whole population (i.e. colonize it) before drug addition. Thus, the overall colonization time becomes smaller than $t_{\textrm{add}}$ and tends toward a new plateau. Note that in this regime, colonization can occur before drug addition through migrations or new mutations, followed by fixations. The time it takes for a locally successful mutant to appear in the slowest deme can be expressed as $\langle t_{af\,S}(N^{*},D)\rangle=\langle t_{af\,W}(N^{*})\rangle\sum_{i=1}^{D}1/i$ (see Supplementary Appendix Section 1.3). It is beyond this time that we expect a plateau even for small migration rates. This is indeed what is observed in Fig 5B.
Extensions to more complex spatial structures and heterogeneous inoculum
Rescue by resistance is also facilitated by spatial structures beyond the clique.
So far, we considered a minimal model of spatially structured population, known as the clique or the island model, where all demes are equivalent and connected to one another by identical migration rates (see “Models and methods”). We showed that spatial structure facilitates evolutionary rescue through the local fixation of resistant mutants in demes. This mechanism can also exist in different spatial structures, where demes are placed on graphs other than a clique.